determine the momentum of a 0.45 caliber bullet (m= 0.162 kg) leaving the muzzle of a gun at 850 m/s

determine the momentum of a 100-kg professional fullback running across the line at 9.0 m/s

a bicycle has a momentum of 20 kg*m/s. what momentum would the bicycle have if it had the same mass and was moving with twice the speed?

a bicycle has a momentum of 20 kg*m/s. what momentum would the bicycle have if it had one-half the mass and was moving with twice the speed?

a bicycle has a momentum of 20 kg*m/s. what momentum would the bicycle have if it had the same mass and was moving with one half the speed?
Answer #1

a)

Momentum is given by

P=mV =0.162*850=137.7 Kg-m/s

b)

P=100*9=900 Kg-m/s

c)

=>P2/P1=V2/V1

P2/20 =(2V/V)

P2=40 Kg-m/s

d)

P2/P1=(m2/m1)(V2/V1)

P2/20 =(m/2/m)(2V/V)=1

P2=20 Kg-m/s

e)

P2/P1=V2/V1

P2/20 =(V/2/V)=1/2

P2=10 Kg-m/s